# EPW varieties as moduli spaces on ordinary GM surfaces and special GM threefolds

### v2

@@ 6.3. Double EPW surfaces as moduli spaces
A generic special Gushel–Mukai threefold $X$ does not have a special Gushel–Mukai threefold as its period dual according to [31, Lemma 3.8]. On the other hand, one can instead find an ordinary Gushel–Mukai threefold as its period dual by [31, Lemma 3.8]. So one can obtain
@@ 6.4. Description of the special period partners
One notices that the Kuznetsov component used in [29] is different from ${\textbf{{Ku}}}(X)$ but equivalent. The relation between the distinguished stable object $G_{X}$ and the characterizing object in [29, Theorem 1.1] is explained in [28, Section 5]. The statement about the stable objects in the Brill–Noether locus follows from the argument for [29, Theorem 1.1]. ∎
The representatives for other singular points in the moduli space $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})$ have been determined in [33, Corollary 6.11]. Each of them is given by a conic in the ramification locus of the covering involution on $X$ and has the form $\mathscr{I}_{C}[1]$. In particular, one has
@@
The object $K_{x}$ is in the locus $\operatorname{BN}_{X}(\mathscr{I}_{C}[1])$ if and only if $x\notin C$. So one has a birational map $X_{C}\dashrightarrow X$ whose exceptional locus on $X$ is the conic $C$. It seems that $X_{C}$ is exactly the elementary transformation of $X$ along the conic $C$ (see [27, Section 4.4]).
@@ Acknowledgments
The first author would like to thank his advisors Laura Pertusi and Paolo Stellari for their constant supports, and thank Arend Bayer and Jieao Song for helpful conversations. The authors would like to thank Kieran O’Grady and Sasha Kuznetsov for answering their questions.

### v3

@@ 6.3. Double EPW surfaces as moduli spaces
A generic special Gushel–Mukai threefold $X$ does not have a special Gushel–Mukai threefold as its period dual according to [34, Lemma 3.8]. On the other hand, one can instead find an ordinary Gushel–Mukai threefold as its period dual by [34, Lemma 3.8]. So one can obtain
@@ 6.4. Description of the special period partners
One notices that the Kuznetsov component used in [32] is different from ${\textbf{{Ku}}}(X)$ but equivalent. The relation between the distinguished stable object $G_{X}$ and the characterizing object in [32, Theorem 1.1] is explained in [31, Section 5]. The statement about the stable objects in the Brill–Noether locus follows from the argument for [32, Theorem 1.1]. ∎
The representatives for other singular points in the moduli space $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})$ have been determined in [37, Corollary 6.11]. Each of them is given by a conic in the ramification locus of the covering involution on $X$ and has the form $\mathscr{I}_{C}[1]$. In particular, one has
@@
One can see that $\operatorname{BN}_{X}(\mathscr{I}_{C}[1])$ does not intersect with $\operatorname{BN}_{X}(G_{X})$ in $\textsf{{M}}_{\sigma_{X}}(\kappa_{1}+2\kappa_{2})$ as
$\dim\operatorname{Hom}({\textup{{{P}}}}\mathscr{O}_{x},\mathscr{I}_{C}[3])=\dim\operatorname{Hom}(\mathscr{O}_{x},\mathscr{I}_{C}[3])=\dim\operatorname{Hom}(\mathscr{I}_{C},\mathscr{O}_{x})<3$
while it is expected that $X_{C}$ is the elementary transformation of $X$ along the conic CC. [^2] The readers can find the construction of elementary transformations in [30, Section 4.4].
Appendix A On the period of double EPW sextics
Here, we will introduce the period domain $\mathbb{D}_{\Lambda}^{\textup{BB}}$ of double EPW sextics and the involution on $\mathbb{D}_{\Lambda}^{\textup{BB}}$ induced by sending $A$ to $A^{\perp}$. In particular, we will show Proposition A.3.
A.1. The period domain and the moduli space
The exposition here is copied from [44] and one can find a more detailed treatment for lattice theory in [27, Chapter 14].
Let $U_{1},U_{2},U_{3}$ be three copies of the hyperbolic plane $U$, let $E^{8}(-1)$ be the unique unimodular negative definite even lattice of rank $8$, and let $\mathbb{Z}(-2)$ the rank one lattice with a generator of square (2)(-2). Then one defines a lattice
$\tilde{\Lambda}:=U_{1}\oplus U_{2}\oplus U_{3}\oplus E_{8}(-1)^{2}\oplus(-2)\cong U^{3}\oplus E_{8}(-1)^{2}\oplus\mathbb{Z}(-2)$
[…]
$\Lambda:=v_{1}^{\perp}\cong U^{2}\oplus E_{8}(-1)^{2}\oplus\mathbb{Z}(-2)^{2}$
is isometric to the primitive part of $H^{2}(X,\mathbb{Z})$ when $X$ has a polarization of square $2$.
The isometry group $O(\Lambda)$ acts naturally on the discriminant group $A_{\Lambda}=\Lambda^{\vee}/\Lambda$, and one defines the stable orthogonal group $\tilde{O}(\Lambda)$ to be the kernel of $O(\Lambda)\rightarrow O(A_{\Lambda})$.
Now, one can define the fundamental domain
$\Omega_{\Lambda}:=\{[\omega]\in\mathbb{P}(\Lambda_{\mathbb{C}})\,|\,q(\omega,\omega)_{\mathbb{C}}=0\textup{ and }q(\omega,\bar{\omega})_{\mathbb{C}}>0\}$
associated with the lattice $\Lambda$, and the period domain $\mathbb{D}_{\Lambda}:=\Omega_{\Lambda}/\tilde{O}(\Lambda)$. It is known that $\mathbb{D}_{\Lambda}$ is a quasi-projective variety [2] and one denotes $\mathbb{D}_{\Lambda}^{\textup{BB}}$ by its Baily–Borel compactification.
Then one considers the symplectic Grassmannian $\operatorname{LG}(\mathsf{\Lambda}^{3}V_{6})$ which parameterizes Lagrangian subspaces with respect to the symplectic form given by wedge-product. There are two prime divisors $\Sigma$ and $\nabla$ in $\operatorname{LG}(\mathsf{\Lambda}^{3}V_{6})$ due to [44]. The divisor $\Sigma$ is formed by the Lagrangian subspaces which contain decomposable vectors, and the divisor $\nabla$ consists of the Lagrangian subspaces $A$ such that $Y^{\geq 3}_{A}\neq\varnothing$. The moduli space of double EPW sextics is defined as the GIT quotient
and the divisor $\Sigma$ descends to a closed subspace $\mathfrak{R}$ inside $\mathfrak{M}_{EPW}$.
A.2. The period map and involutions
There exists a rational map
$\mathfrak{p}\colon\mathfrak{M}_{EPW}\dashrightarrow\mathbb{D}_{\Lambda}^{\textup{BB}}$
defined by the hyperkähler period map on the open locus in $\mathfrak{M}_{EPW}$ where double EPW sextic is a hyperkähler fourfold. In fact, $\mathfrak{p}$ is regular on $\mathfrak{M}_{EPW}$-$\mathfrak{R}$ by [44, Theorem 1.2].
In a precise way, suppose that the double EPW sextic $X=\tilde{Y}_{A}^{\geq 1}$ is hyperkähler, then it has a natural polarization $h$ of square $2$ by [41]. So there exists an isometry
$\psi\colon H^{2}(X,\mathbb{Z})\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\tilde{\Lambda},\quad\psi(h)=v_{1}$
[…]
Theorem A.1 ([44, Theorem 1.3]).
The restriction of $\mathfrak{p}$ on $\mathfrak{M}_{EPW}$-$\mathfrak{R}$ goes into the open subset $\mathbb{D}_{\Lambda}-\mathbb{S}_{4}$, where $\mathbb{S}_{4}:=\rho(\Omega_{\Lambda}\cap(e_{1}+e_{2})^{\perp})$ for the quotient $\rho\colon\Omega_{\Lambda}\rightarrow\mathbb{D}_{\Lambda}$.
The mapping $A\mapsto A^{\perp}$ induces an involution $\delta$ on $\mathfrak{M}_{EPW}$, and one can also define it at the level of period domains. Here let us choose a generator $e_{1}$ of $v_{1}^{\perp}\cap U_{1}$ and $e_{2}$ be a generator of the second $\mathbb{Z}(-2)$ in $\Lambda$, then one has an involution $\iota\in O(\Lambda)$ characterized by
$\iota(e_{1})=e_{2},\quad\iota(e_{2})=e_{1},\quad\iota|_{\{e_{1},e_{2}\}^{\perp}}=\operatorname{id}_{\{e_{1},e_{2}\}^{\perp}}$
which descends to a non-trivial involution $\bar{\iota}$ on $\mathbb{D}_{\Lambda}$ and then extends to $\bar{\iota}\colon\mathbb{D}_{\Lambda}^{\textup{BB}}\rightarrow\mathbb{D}_{\Lambda}^{\textup{BB}}$.
Theorem A.2 (O’Grady [42, 44]).
One has $\mathfrak{p}\circ\delta=\bar{\iota}\circ\mathfrak{p}$ for the involution $\delta$ on $\mathfrak{M}_{EPW}$.
Here only the involution $\bar{\iota}\colon\mathbb{D}_{\Lambda}\rightarrow\mathbb{D}_{\Lambda}$ is relevant. One chooses an element $e_{3}\in U_{2}\subset\tilde{\Lambda}$ such that $e_{3}^{2}=-2$, then according to [44, Proposition 4.3] the $\tilde{O}(\Lambda)$-orbits of $e_{1},e_{2},e_{3},e_{1}+e_{2}$ are the four distinguished points in $\mathbb{D}_{\Lambda}$ with negative squares. So one can see that $\bar{\iota}$ is in fact the reflection with respect to the $\tilde{O}(\Lambda)$-orbit of $e_{1}+e_{2}$.
Proposition A.3.
The involution $\delta$ on $\mathfrak{M}_{EPW}$ does not fix any point in $\mathfrak{M}_{EPW}$-$\mathfrak{R}$.
Proof.
Suppose that $\bar{\iota}([\omega])=[\omega]$ for some element in $\mathbb{D}_{\Lambda}$, then $\iota(\omega)=c\omega$ for some $c\neq 0$. It follows that $c=\pm 1$, as $\iota$ is also an involution on $\Lambda_{\mathbb{C}}$. Since $\omega^{2}=0$ and $\iota$ is a reflection with respect to the element $e_{1}-e_{2}$, one can exclude the case c=1c=-1. So one has $\omega\in\Omega_{\Lambda}\cap(e_{1}-e_{2})^{\perp}$ which follows that the class $\omega]$ is in $\mathbb{S}_{4}$ as the elements $e_{1}\pm e_{2}$ are in the same orbit.
Then one concludes by applying Theorem A.1 and Theorem A.2. ∎
@@ Acknowledgments
The first author would like to thank his advisors Laura Pertusi and Paolo Stellari for their constant supports, and thank Arend Bayer and Jieao Song for helpful conversations. The authors would like to thank Kieran O’Grady and Sasha Kuznetsov for answering their questions. The authors also want to thank the anonymous referees for beneficial comments.